Suppose a consumer consuming two commodities x and y has the following utility function u=x⁰⁴y⁰⁶. If price of good x and good y are birr 4 and birr 6 respectively, and income constraint is birr 100.a.find marginal rate of substitution of x for y and y for x. B.find the quantities of x and y which maximize utility. C.show how a rise in income to birr 200 will affect the quantities of x and y at equilibrium.
1
Expert's answer
2021-10-25T17:40:40-0400

To maximize utility: MRSxy "=" "\frac{MUx}{MUy} = \frac{Px}{Py}"


Derive MUx:


U = 10X0.4Y0.6


MUx = "\frac{\partial U} {\partial X}" = 0.4X-0.6Y0.4


Derive MUy:


MUy = "\frac{\partial U} {\partial Y}" = 0.6X0.4Y-0.4


MRSxy = "\frac{MUx}{MUy}" = "\frac{0.4X^{-0.6}Y^{0.4} }{0.6X^{0.4}Y^{-0.4} }"


MRSxy = "\frac{Px}{Py}"


"\frac{0.2Y^{0.8} }{0.3X }" = "\frac{0.2 }{0.3 }"


Y = X1.25


Budget constraint: I = PxX + PyY


50 = 2X + 3Y


Substitute the value of X in budget constraint to derive Y:

50 = 2X + 3(X1.25)


50 = 2X + 3X1.25


X = 7


Y = X1.25 = 71.25 = 12


Y = 12


The quantities of X and Y which maximize utility (Uxy) = (7, 12)

New budget constraint: 100 = 2X + 3Y


Y = X1.25


Substitute the value of X in budget constraint to derive Y:


100 = 2X + 3(X1.25)


100 = 2X + 3X1.25

X = 13


Y = X1.25 = 131.25 = 25


Y = 25


The new quantities of X and Y which maximize utility (Uxy) = (13, 25)